OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.55 2.25 2.95 3.65 Which equation gives the cost, y, in terms of the number of pages, x? y = 0.014x + 49.15 y = 0.014x + 1.55 y = 0.014x + 0.85 y = 58.8x + 0.85

OpenStudy (anonymous):

@musicdramalife00 ?

OpenStudy (anonymous):

first find the slope

OpenStudy (anonymous):

(3.65 -2.95 )/ (200-150) =m

OpenStudy (anonymous):

mkmk and then you use like the point-slop form you can just take a random point hmmmm lets say 200, 3.65 and m= 0.014 plug it into the equation and solve

OpenStudy (anonymous):

soooooo like y-3.65= 0.014 (x-200) and you distribute 0.013x-2.8=y-3.65

OpenStudy (anonymous):

add 3.65 on both sides

OpenStudy (anonymous):

y=0.014x +0.85